OpenStudy (anonymous):

Uranium-234 is unstable and undergoes two subsequent alpha emissions. What is the resulting nuclide from this transformation?

OpenStudy (anonymous):

The resulting nuclide is Uranium-226. (I think.)

OpenStudy (frostbite):

alpha decay is the emission of 4/2 He. so it can't be uranium it self.

OpenStudy (frostbite):

Look at the periodic table and go 4 elements back from uranium.

OpenStudy (anonymous):

Alright! Thank you again! :D Goodnight Frostbite! :P

OpenStudy (frostbite):

But the nuclei number is correct btw :) and good night,

OpenStudy (anonymous):

I got Yb? hahaha what the heck?

OpenStudy (frostbite):

Yb? sure? I want it to be Radium as it has the atomic number 88 where uranium has the atomic number 92. 92-2*2=92-4=88 (can't write as you might can see)

OpenStudy (anonymous):

Oh my goodness, I forgot the the periodic table is cut into sections! Ahh I see now. :)

OpenStudy (frostbite):

Perfect now I can faint in my bed with good conscience.

OpenStudy (anonymous):

Goodnight. hahahaha :P